OpenStudy (anonymous):

Solve for the variables. y = - 3 – z - 2 - 2 + y = - 3 + z Question 2 options: a) y = -2; z = -3 b) y = -3; z = 2 c) y = -3; z = -2 d) None of these

OpenStudy (anonymous):

I tried substitution, but I did not get any of these. Please help.

OpenStudy (johnweldon1993):

Lets try it again y = - 3 – z - 2 - 2 + y = - 3 + z replace 'y' with the first equation -2 -3 - z - 2 = -3 + z -5 - z - 2 = -3 + z -7 - z = -3 + z -4 - z = z -4 = 2z z = -2 Now use z to solve for 'y' y = - 3 – (-2) - 2 y = -3 + 2 - 2 y = -3 so I get z = -2 y = -3

OpenStudy (anonymous):

Thank you!!

OpenStudy (johnweldon1993):

No problem! :)